# Chasing the Tail: Effective Rubric-based Reward Modeling for Large Language Model Post-Training

### v1

@@ Abstract
Reinforcement fine-tuning (RFT) often suffers from reward over-optimization, where a policy model hacks the reward signals to achieve high scores while producing low-quality outputs. Our theoretical analysis shows that the key lies in reward misspecification at the high-reward tail: the inability to reliably distinguish excellent responses from merely great ones. This motivate us to focus on the high-reward region. However, such tail examples are scarce under the base LLM. While off-policy exemplars (e.g. from stronger models or rewrites) are easier to obtain, naively training on them yields a misspecified reward for the policy we aim to align. To address this, we study rubric-based rewards. By design, rubrics can leverage off-policy examples while remaining insensitive to their artifacts. To elicit rubrics that capture the high-reward tail, we highlight the importance of distinguishing among great and diverse responses, and introduce a workflow to implement this idea. We empirically demonstrate that rubric-based rewards substantially mitigate reward over-optimization and deliver effective LLM post-training improvements. [^1] Our code can be accessed at https://github.com/Jun-Kai-Zhang/rubrics.git.
[…]
@@ 1 Introduction
In this paper, we are interested in how to produce reward models that are effective when used for LLM post-training. A reward model is a function that takes a prompt and a response and produces a score quantifying how good that response is for the prompt. In post-training, we then align a language model to the reward by a reinforcement-learning type procedure. The fundamental challenge here is that, in many settings, it is nearly inevitable that the reward model will be an imperfect proxy for the behavior that we are actually trying to induce. In particular, this means that as we run post-training, it will increasingly be the case that the LLM is aligned to the idiosyncratic misspecification of the reward rather than the true signal that we are trying to extract. In this paper, we are interested in mitigating this effect.
Given that some misspecification is inevitable, what should we focus on when defining a reward model? The basic setup of post-training aims to induce the good behavior encoded by the reward while minimally shifting other aspects of the base LLM. Mathematically, this can be formalized as looking for post-training procedures that move along the Pareto frontier of KL divergence from the base model vs win-rate (as judged by the reward) against the base model. We begin by theoretically demonstrating that, for such Pareto-optimal procedures, the effect of reward misspecification is dominated by errors in the high-reward region. In other words, what really matters for post-training is the ability to accurately distinguish between the very good responses.
Then, we know that we want to focus our reward modeling on the high-reward region of examples. The basic challenge here is that actually producing high-reward examples to train a reward model on is hard. If we simply sample responses from the base LLM itself, then it is extremely sample inefficient to get the necessary examples (because we are trying to get elements in a low-probability tail). On the other hand, if we use an off-policy procedure—e.g., drawing samples from a stronger LLM, or producing good examples with extensive thinking or rewrites—we can get high-reward examples, but naively training a reward model on them may learn superficial features instead of eliciting real capabilities (see Appendix˜D).
@@ 2 Preliminaries
[…]
@@
[…]
where is a hyperparameter to control fine-tuned model’s deviation from the reference model, i.e.,
[…]
As demonstrated in Rafailov et al. (2023), the solution to (2.1) is
[…]
@@
Reinforcement learning from rubrics-based reward (RLRR) (Gunjal et al., 2025; Viswanathan et al., 2025; Huang et al., 2025b) has emerged as a promising approach for open-ended tasks. The core idea is to associate each prompt $x$ with a rubric—a set of explicit criteria ($c_{i}$) with corresponding weights ($w_{i}$) that collectively define a high-quality response. For instance, given a prompt asking for a likely diagnosis from a patient’s symptoms, the rubric could specify key aspects of a good answer. This might include high-weight criteria for “identifying [correct diagnosis] as the likely diagnosis” and “correctly identifying the condition as a medical emergency,” and a low-weight criterion for “mentioning [typical treatment] for treatment” (See Appendix˜I for a concrete example.)
@@ 3 High-Reward Region Accuracy is Key to Overcoming Reward Over-optimization
To characterize the reward over-optimization phenomenon, we need to study the relationship between the utility (expected reward and win rates), and the KL divergence in (2.2). They can be simplified as follows:
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
The explicit formula in Theorem˜2 indicates that misspecification, i.e., the deviation of $f$ from the identity map, in the high-value region of $r^{\star}$ has dominantly large effects on the utility-KL tradeoff. On one hand, the KL divergence remains invariant to the choice of $f$ and is fixed when the penalty parameter is set. On the other hand, the exponential term imposes increasingly severe penalties on misspecification in the high-reward regime relative to the low-reward regime. This highlights the criticality of accuracy in the high-reward region for achieving a favorable balance between utility and KL divergence.
@@
Figure˜2(a) plots KL divergence versus win rate across misspecification patterns and yields two key observations: (i) when the proxy is inaccurate in the high-reward region, performance may look acceptable at small KL but the win rate collapses as KL grows (this is similar to the reward over-optimization behavior in Gao et al. (2023)); and (ii) if the proxy correctly ranks just a small top proportion of responses (e.g., $10\%$), even while misgrading the remaining majority, the win rate rapidly approaches the optimal curve at moderate KL. Separately, Figure˜2(b) varies the fraction $c$ of correctly ranked top responses and traces the corresponding lower envelope of achievable win rates, showing that this envelope is already near-optimal once a sufficiently large top proportion is correctly identified and ordered (e.g., $40\%$). Together, we reach our central theoretical findings:
@@ 4 Principles for Constructing Rubrics
Based on the results of the previous section, we construct a reward model focusing on the high-value region. The problem then is getting training examples that are in this high-reward region. By definition, these are samples that are rare under the base LLM policy! This essentially forces us to use off-policy data to define the reward model. Now, rubric-based rewards have emerged as an approach for using off-policy data to define rewards. The basic idea of rubric-based reward models is to explicitly restrict the reward to only care about aspects of the solution that are relevant to its quality, thereby mitigating the effect of the off-policy data. However, the restrictive nature of the rubrics is a double-edged sword. The same structure that limits the effect of off-policyness may also limit their ability to distinguish between solutions that are excellent and those that are merely great (they can easily end up in a tie). In this section, we consider how to construct rubrics that are focused on accuracy in the high-reward region.
@@
Table 1: RL experimental results on three datasets across two domains. Both principles lead to consistent improvements.
Method Generalist Domain Health Domain
Filtered Set LMArena Medical-o1
Win-Rate % Win-Rate % Win-Rate % HealthBench
Base Policy 5.2 4.1 10.8 $0.1721$
SFT 35.9 29.6 25.8 $0.2999$
Initial, Prompt only 31.3 29.7 21.7 $0.3004$
1 Good Pair 33.5 32.8 22.4 $0.2912$
1 Great Pair 36.8 33.1 26.5 0.3163
4 Great Pairs 38.7 34.7 31.4 $0.3348$
4 Great & Diverse Pairs 39.7 35.1 34.4 0.3513
@@ 4.1 Methodology
While a single RTD step sharpens the rubric, repeated application over a larger candidate pool yields systematic improvements. Starting with all off-policy responses for a prompt, each iteration scores the candidates under the current rubric, selects the top two responses, and refines the rubric using RTD. This workflow concentrates rubric discovery on the performance frontier, extracting the most informative distinctions from the best available responses with only a small number of comparisons (see Algorithm˜1 and Figure˜3).
@@ 4.2 Experimental Setup
We conduct studies in two distinct domains: general-purpose and professional (health). More concretely, we set up the experiments as follows:
@@
We employ GPT-4.1 as the rubric proposer, and prompt it to get the initial rubrics. The training datasets consist of two generalist prompt collection (LMArena (Chiang et al., 2024) and a manually filtered set of natural prompts, detailed in Appendix˜G) and one technical, domain-specific prompt set (medical-o1-reasoning-SFT (Chen et al., 2024)). Each dataset contains $5000$ prompts for training and an additional $1000$ prompts for in-domain evaluation. The base model for post-training is Qwen3-8b-Base (Yang et al., 2025), which has instruction-following capabilities. We adopt GRPO (Shao et al., 2024) as the RFT algorithm and use a standard set of hyperparameters, detailed in Table 4. For the reward computation, we leverage GPT-4.1-mini as a rubric verifier and calculate the final reward as the weighted sum of satisfied rubric criteria, normalized by the total weight. All prompts used in the experiments are presented in Appendix˜B.
@@
To validate Principle 1, we compare rubrics refined using (i) candidate pairs from a great model versus (ii) candidate pairs from a good model (Gemini 2.5 Pro and Gemini-2.5-Flash-Lite, respectively (Comanici et al., 2025)). To validate Principle 2, we enlarge the pool by sampling 16 responses per prompt, from a broader set of excellent models, ensuring greater diversity (see Appendix˜F for the full list). This setup allows us to test whether rubric refinement benefits from better and more diverse candidate responses.
Evaluations.
Final performance is evaluated by head-to-head comparison against Qwen3-8B, a strong hybrid thinking model, using the held-out set of $1000$ prompts from each dataset. We compute the win-rate based on judgments from our GPT-4.1 judge , which was prompted to act as an impartial evaluator (see Appendix E for the detail and justifications). For health domain, we additionally evaluate models on the HealthBench (Arora et al., 2025).
@@ 5.1 RL improves with better and more diverse responses
We first evaluate downstream RL performance to test whether the proposed principles indeed improve rubrics. Table˜1 shows two clear trends. First, rubrics refined with great pairs outperform those refined with good pairs, validating Principle 1. Second, iterative refinement with multiple diverse great pairs yields further gains, validating Principle 2.
Figure 4: Refinement with great and diverse responses mitigates reward over-optimization. Training rewards $r$ rise similarly across settings, but only models trained with iteratively refined, diverse rubrics sustain higher win-rates (a proxy for ground-truth reward $r^{*}$) and delay the collapse that signals reward over-optimization.
Beyond improving average performance, refinement with better and more diverse responses also mitigates reward over-optimization. Figure˜4 shows training dynamics on the health domain when RL is run for extended steps. Models trained on initial rubrics, or rubrics refined with only a single pair, peak early and then suffer a rapid decline in win rate after about 60 steps—an indicator of reward over-optimization. In contrast, models trained with iteratively refined, diverse rubrics sustain higher win-rates for much longer, with over-optimization not appearing until after roughly 160 steps. This pattern indicates that refining rubrics with great and diverse responses corrects inaccuracies in the high-reward region, thereby delaying the onset of over-optimization. Together, these results confirm our central hypothesis that rubrics can be constructed to mitigate reward over-optimization.

### v2

@@ Abstract
Reinforcement fine-tuning (RFT) often suffers from reward over-optimization, where a policy model hacks the reward signals to achieve high scores while producing low-quality outputs. Our theoretical analysis shows that the key lies in reward misspecification at the high-reward tail: the inability to reliably distinguish excellent responses from merely great ones. This motivate us to focus on the high-reward region. However, such tail examples are scarce under the base LLM. While off-policy exemplars (e.g. from stronger models or rewrites) are easier to obtain, naively training on them yields a misspecified reward for the policy we aim to align. To address this, we study rubric-based rewards. By design, rubrics can leverage off-policy examples while remaining insensitive to their artifacts. To elicit rubrics that capture the high-reward tail, we highlight the importance of distinguishing among great and diverse responses, and introduce a workflow to implement this idea. We empirically demonstrate that rubric-based rewards substantially mitigate reward over-optimization and deliver effective LLM post-training improvements.***Our code can be accessed at https://github.com/Jun-Kai-Zhang/rubrics.git.
11footnotetext: Equal contribution.22footnotetext: Work done during internship at Scale AI.
@@ 1 Introduction
In this paper, we are interested in how to produce reward models that are effective when used for LLM post-training. A reward model is a function that takes a prompt and a response and produces a score quantifying how good that response is for the prompt. In post-training, we then align a language model to the reward by a reinforcement learning type procedure. The fundamental challenge here is that, in many settings, it is nearly inevitable that the reward model will be an imperfect proxy for the behavior that we are actually trying to induce. In particular, this means that as we run post-training, it will increasingly be the case that the LLM is aligned to the idiosyncratic misspecification of the reward rather than the true signal that we are trying to extract. In this paper, we are interested in mitigating this effect.
Given that some misspecification is inevitable, what should we focus on when defining a reward model? The basic setup of post-training aims to induce the good behavior encoded by the reward while minimally shifting other aspects of the base LLM. Mathematically, this can be formalized as looking for post-training procedures that move along the Pareto frontier of KL divergence from the base model vs win rate (as judged by the reward) against the base model. We begin by theoretically demonstrating that, for such Pareto-optimal procedures, the effect of reward misspecification is dominated by errors in the high-reward region. In other words, what really matters for post-training is the ability to accurately distinguish between the very good responses.
Then, we know that we want to focus our reward modeling on the high-reward region of examples. The basic challenge here is that actually producing high-reward examples to train a reward model on is hard. If we simply sample responses from the base LLM itself, then it is extremely sample inefficient to get the necessary examples (because we are trying to get elements in a low-probability tail). On the other hand, if we use an off-policy procedure—e.g., drawing samples from a stronger LLM, or producing good examples with extensive thinking or rewrites—we can get high-reward examples, but naively training a reward model on them may learn superficial features instead of eliciting real capabilities (see Appendix D).
@@ 2 Preliminaries
We use $\pi$ to denote a large language model (LLM) and $\pi_{0}$ to denote the reference language model (usually the starting point of RL). Given a prompt $x$, a response $y$ is sampled from the conditional distribution $\pi(\cdot\mid x)$. A reward model $r(\cdot,\cdot)$ is utilized to assess the quality of a prompt-response pair. We use $r^{\star}$ to represent the gold reward model (inaccessible in practice) and $r$ to represent the proxy reward applied in practice.
@@
$\max_{\pi}\mathbb{E}_{x\sim D,~y\sim\pi(\cdot\mid x)}\left[r(x,y)\right]-\beta\mathbb{D}_{\mathrm{KL}}\left[\pi(y\mid x)\|\pi_{0}(y\mid x)\right],$ (1)
where $\beta$ is a hyperparameter to control fine-tuned model’s deviation from the reference model, i.e.,
$\mathbb{D}_{\mathrm{KL}}\left[\pi(y\mid x)\|\pi_{0}(y\mid x)\right]=\mathbb{E}_{x\sim D,y\sim\pi(\cdot\mid x)}\left[\log\frac{\pi(y\mid x)}{\pi_{0}(y\mid x)}\right].$
As demonstrated in Rafailov et al. (2023), the solution to (1) is
$\pi_{r}(y\mid x)\propto\pi_{0}(y\mid x)\exp\{r(x,y)/\beta\}.$ (2)
@@
Reinforcement learning from rubrics-based reward (RLRR) (Gunjal et al., 2025; Viswanathan et al., 2025; Huang et al., 2025b) has emerged as a promising approach for open-ended tasks. The core idea is to associate each prompt $x$ with a rubric—a set of explicit criteria ($c_{i}$) with corresponding weights ($w_{i}$) that collectively define a high-quality response. For instance, given a prompt asking for a likely diagnosis from a patient’s symptoms, the rubric could specify key aspects of a good answer. This might include high-weight criteria for “identifying [correct diagnosis] as the likely diagnosis” and “correctly identifying the condition as a medical emergency,” and a low-weight criterion for “mentioning [typical treatment] for treatment” (See Appendix I for a concrete example.)
@@ 3 High-Reward Region Accuracy is Key to Overcoming Reward Over-optimization
To characterize the reward over-optimization phenomenon, we need to study the relationship between the utility (expected reward and win rates), and the KL divergence in (2). They can be simplified as follows:
@@
Define $R^{x}_{0}=r^{\star}(x,Y_{0})$ with $Y_{0}\sim\pi_{0}(\cdot\mid x)$ and $F^{x}_{0}$ as its cumulative distribution function. The RFT solution (2) has:
@@
    Expected reward: $\mathbb{E}_{x\sim D,~y\sim\pi_{r}(\cdot\mid x)}\left[r^{\star}(x,y)\right]=\mathbb{E}_{x\sim D}\left[\frac{\mathbb{E}\left[R^{x}_{0}~e^{f(R^{x}_{0})/\beta}\right]}{\mathbb{E}\left[e^{f(R^{x}_{0})/\beta}\right]}\right]$,
@@
    Win Rate: $\mathbb{E}_{x\sim D,~y\sim\pi_{r}(\cdot\mid x)}\left[F_{0}^{x}\left(r^{\star}(x,y)\right)\right]=\mathbb{E}_{x\sim D}\left[\frac{\mathbb{E}\left[F^{x}_{0}(R^{x}_{0})~e^{f(R^{x}_{0})/\beta}\right]}{\mathbb{E}\left[e^{f(R^{x}_{0})/\beta}\right]}\right]$,
@@
    KL divergence: $\mathbb{D}_{\mathrm{KL}}\left[\pi_{r}(y\mid x)\|\pi_{0}(y\mid x)\right]=\mathbb{E}_{x\sim D}\left[\frac{\mathbb{E}\left[f(R^{x}_{0})~e^{f(R^{x}_{0})/\beta}/{\beta}\right]}{\mathbb{E}\left[e^{f(R^{x}_{0})/\beta}\right]}-\log\mathbb{E}\left[e^{f(R^{x}_{0})/\beta}\right]\right]$
@@
    $\mathbb{D}_{\mathrm{KL}}\left[\pi_{r}(y\mid x)\|\pi_{0}(y\mid x)\right]=\frac{(1/\beta-1)e^{1/\beta}+1}{e^{1/\beta}-1}-\log\beta-\log(e^{1/\beta}-1).$
@@
    Expected reward (or win rate) of $\pi_{r}$ is $\frac{\int_{0}^{1}f^{-1}(u)e^{u/\beta}\mathrm{d}u}{\beta\left(e^{1/\beta}-1\right)}$. [Proof].
The explicit formula in Theorem 1 indicates that misspecification, i.e., the deviation of $f$ from the identity map, in the high-value region of $r^{\star}$ has dominantly large effects on the utility-KL tradeoff. On one hand, the KL divergence remains invariant to the choice of $f$ and is fixed when the penalty parameter $\beta$ is set. On the other hand, the exponential term imposes increasingly severe penalties on misspecification in the high-reward regime relative to the low-reward regime. This highlights the criticality of accuracy in the high-reward region for achieving a favorable balance between utility and KL divergence.
@@
Figure 2(a) plots KL divergence versus win rate across misspecification patterns and yields two key observations: (i) when the proxy is inaccurate in the high-reward region, performance may look acceptable at small KL but the win rate collapses as KL grows (this is similar to the reward over-optimization behavior in Gao et al. (2023)); and (ii) if the proxy correctly ranks just a small top proportion of responses (e.g., $10\%$), even while misgrading the remaining majority, the win rate rapidly approaches the optimal curve at moderate KL. Separately, Figure 2(b) varies the fraction $c$ of correctly ranked top responses and traces the corresponding lower envelope of achievable win rates, showing that this envelope is already near-optimal once a sufficiently large top proportion is correctly identified and ordered (e.g., $40\%$). Together, we reach our central theoretical findings:
@@ 4 Principles for Constructing Rubrics
Based on the results of the previous section, we construct a reward model focusing on the high-value region. The problem then is getting training examples that are in this high-reward region. By definition, these are samples that are rare under the base LLM policy! This essentially forces us to use off-policy data to define the reward model. Now, rubric-based rewards have emerged as an approach for using off-policy data to define rewards. The basic idea of rubric-based reward models is to explicitly restrict the reward to only care about aspects of the solution that are relevant to its quality, thereby mitigating the side effect of the off-policy data. However, the restrictive nature of the rubrics is a double-edged sword. The same structure that limits the effect of off-policyness may also limit their ability to distinguish between solutions that are excellent and those that are merely great (they can easily end up in a tie). In this section, we consider how to construct rubrics that are focused on accuracy in the high-reward region.
@@
Table 1: RL experimental results across three domains. The base policy model is Qwen3-8B-Base, and the win rate is evaluated against Qwen3-8B. The results show two clear trends: refining rubrics by differentiating two great responses outperforms differentiating two good responses, and differentiating among a diverse set of great responses further improves performance.
Method Generalist Domain Health Domain Finance Domain
Filtered Set LMArena Medical-o1 Finance
Win% Win% Win% Score Win % Score
Base Policy 5.2 4.1 10.8 $0.1721$ 5.84 0.1738
SFT 35.9 29.6 25.8 $0.2999$ 26.04 0.2218
Initial, Prompt only 31.3 29.7 21.7 $0.3004$ 37.23 0.2693
1 good Pair 33.5 32.8 22.4 $0.2912$ 39.07 0.2694
1 great Pair 36.8 42.24 26.5 0.3163 42.24 0.2838
4 great Pairs 38.7 34.7 31.4 $0.3348$ 48.91 0.2961
4 great & Diverse Pairs 39.7 35.1 34.4 0.3513 49.58 0.3018
@@ 4.1 Methodology
While a single RTD step sharpens the rubric, repeated application over a larger candidate pool yields systematic improvements. Starting with all off-policy responses for a prompt, each iteration scores the candidates under the current rubric, selects the top two responses, and refines the rubric using RTD. This workflow concentrates rubric discovery on the performance frontier, extracting the most informative distinctions from the best available responses with only a small number of comparisons (see Algorithm 1 and Figure 3).
@@ 4.2 Experimental Setup
Our experimental goals are twofold. First, we examine how leveraging off-policy responses can alleviate reward over-optimization. Second, we assess the efficacy of these methods in enhancing LLM capabilities. Our experiments span three domains: general-purpose, healthcare, and finance. For the first goal, to isolate the effect of rubric refinement strategies, we adopt the synthetic oracle setting proposed by the seminal work Gao et al. (2023). In this framework, a strong LLM acts as a proxy for the “gold-standard” preference source. This oracle model serves a dual purpose: it generates the rubrics data used for reward modeling and serves as the final judge for performance evaluation. By unifying the annotation and evaluation sources, this design eliminates confounding factors arising from preference mismatch between the training data and the judge. This allows us to clearly assess how effectively different rubric refinement strategies mitigate reward over-optimization. For the second goal, we evaluate the model performance on domain-specific objective benchmarks when applicable. Regarding the second goal, we specifically evaluate the model on healthcare and finance tasks. These fields offer established objective professional benchmarks, allowing us to rigorously test improvements in model capabilities. We set up the experiments as follows:
@@
We employ GPT-4.1 as the rubric proposer, prompting it to generate the initial rubrics. The training datasets consist of two generalist prompt collections (LMArena (Chiang et al., 2024) and a manually filtered set of natural prompts, detailed in Appendix G) and two domain-specific prompt sets: for healthcare, we utilize medical-o1-reasoning-SFT (Chen et al., 2024); for the finance domain, we filtered for 1147 high-quality finance prompts in LMArena Team (2025). Each dataset contributes 5000 prompts for training and an additional 1000 prompts for in-domain evaluation, with the exception of the finance dataset, for which we use all available prompts for training and the prompts from the PRBench-Finance (Akyürek et al., 2025) for the evaluation. The base model for post-training is Qwen3-8B-Base (Yang et al., 2025), which has instruction-following capabilities. We adopt GRPO (Shao et al., 2024) as the RFT algorithm and use a standard set of hyperparameters, detailed in Table 4. For the reward computation, we leverage GPT-4.1-mini as a rubric verifier and calculate the final reward as the weighted sum of satisfied rubric criteria, normalized by the total weight. All prompts used in the experiments are presented in Appendix B.
@@
To validate Principle 1, we compare rubrics refined using (i) candidate pairs from a great model versus (ii) candidate pairs from a good model (Gemini 2.5 Pro and Gemini-2.5-Flash-Lite, respectively (Comanici et al., 2025)). To validate Principle 2, we enlarge the pool by sampling 16 responses per prompt, from a broader set of excellent models, ensuring greater diversity (see Appendix F for the full list). This setup allows us to test whether rubric refinement benefits from better and more diverse candidate responses. The supervised fine-tuning (SFT) uses two responses per prompt, sampled from the great model.
Evaluation.
We assess performance across two primary dimensions: (1) alignment with oracle preferences, and (2) domain-specific scores on professional benchmarks. For preference evaluation, we conduct head-to-head comparisons against Qwen3-8B, the strong thinking version of our base policy model, on a held-out set of test prompts. The oracle model GPT-4.1 was prompted to act as an impartial evaluator and select the better response (see Appendix E for the detail). To validate the performance gain in professional domains, we additionally evaluate models on HealthBench (Arora et al., 2025) and PRBench (Akyürek et al., 2025), which provide objective metrics grounded in expert-curated rubrics.
@@ 5.1 RL improves with better and more diverse responses
We first evaluate downstream RL performance to test whether the proposed principles indeed improve rubrics. Table 1 shows two clear trends. First, rubrics refined with great pairs outperform those refined with good pairs, validating Principle 1. Second, iterative refinement with multiple diverse great pairs yields further gains, validating Principle 2.
(a) Healthcare: Win Rate and HealthBench Scores at Different Training Steps
(b) Finance: Win Rate and PRBench-Finance Scores at Different Training Steps
Figure 4: Model performance metrics across training steps for different rubric construction strategies. The figures show the evolution of the Win Rate and respective benchmark scores in the healthcare (4(a)) and finance (4(b)) domains.
Beyond improving average performance, refinement with better and more diverse responses also mitigates reward over-optimization. Figure 4 shows training dynamics on the health and finance domains when RL is run for extended steps. Models trained on initial rubrics, or rubrics refined with only a single pair, peak early and then suffer a rapid decline in win rate and benchmark scores—an indicator of reward over-optimization. In contrast, models trained with iteratively refined, diverse rubrics sustain higher win rates and benchmark scores for much longer, with over-optimization not appearing until late stages. This pattern indicates that refining rubrics with great and diverse responses corrects inaccuracies in the high-reward region, thereby delaying the onset of over-optimization. Together, these results confirm our central hypothesis that rubrics can be constructed to mitigate reward over-optimization.